# On Uncertainty Calibration for Equivariant Functions

### v2

@@ Appendix C Swiss Roll Experiment Details
Define the calibration CE as the term inside the integrand of Equation˜1, $\text{CE}=|\mathrm{Acc}_{p}(h)-p|$. For an i.i.d. set of $n$ samples $\{(x_{1},h(x_{1})),\ldots,(x_{n},h(x_{n}))\}$, we index a given pair by $Z_{i}$. We have that
@@ D.1 Motivation and Implementation of PSF Blurring
A point-spread function (PSF) is an impulse response of an optical system to light. PSFs occur all throughout medical and astronomical imaging. The science case we explore in this work is the distortion of galaxy images. With next generation imagers like JWST and large astronomical surveys like COSMOS-Web (Casey et al., 2023), there are renewed efforts to characterize the effect of the PSF and its effects on downstream scientific analysis (Perrin et al., 2014; Birrer et al., 2021; Jarvis et al., 2021; Michalewicz et al., 2023; Liaudat et al., 2023; Berman et al., 2024; Berman & McCleary, 2024; Feng et al., 2025; Polzin, 2025). Understanding how the PSF harms a model’s ability to identify a galaxy’s morphology class can hint at the effect of the PSF on measured ellipticity moments (Hirata & Seljak, 2003; Mandelbaum et al., 2005), which is a crucial ingredient for maps of large scale structure (McCleary et al., 2015; 2020; Scognamiglio, 2024).
@@ D.2 Training and Evaluating
Our models and training scripts are adapted from Pandya et al. (2025a) and https://github.com/deepskies/SIDDA. The galaxy datasets are initially sourced from https://zenodo.org/records/14583107, and there is a script to produce the datasets with PSF blurring in our GitHub artifact. We compute ECE using the same approximations as in Equations 13 - 15. We summarize the number of parameters for each model in Table˜2 below:
Model CNNCNN $C_{2}$ $C_{4}$ $C_{6}$ $C_{8}$ $C_{10}$ $C_{12}$
@@ Appendix E Chemical Property Experiment Details
We leave further comparison and model development as an opportunity for future work. Other potential baselines could include Equiformer (Liao & Smidt, 2022), EquiformerV2 (Liao et al., 2023), Graphormer (Shi et al., 2022), or Graphormer with data augmentation. The model we use in this work has $36,997,125$ parameters.

### v4

@@ Appendix C Swiss Roll Experiment Details
Define the calibration CE as the term inside the integrand of equation˜2, $\text{CE}=|\mathrm{Acc}_{p}(h)-p|$. For an i.i.d. set of $n$ samples $\{(x_{1},h(x_{1})),\ldots,(x_{n},h(x_{n}))\}$, we index a given pair by $Z_{i}$. We have that
@@ D.1 Motivation and Implementation of PSF Blurring
A point-spread function (PSF) is an impulse response of an optical system to light. PSFs occur all throughout medical and astronomical imaging. The science case we explore in this work is the distortion of galaxy images. With next generation imagers like JWST and large astronomical surveys like COSMOS-Web (Casey et al., 2023), there are renewed efforts to characterize the effect of the PSF and its effects on downstream scientific analysis (Perrin et al., 2014; Birrer et al., 2021; Jarvis et al., 2021; Michalewicz et al., 2023; Liaudat et al., 2023; Berman et al., 2024; Berman and McCleary, 2024; Feng et al., 2025; Polzin, 2025). Understanding how the PSF harms a model’s ability to identify a galaxy’s morphology class can hint at the effect of the PSF on measured ellipticity moments (Hirata and Seljak, 2003; Mandelbaum et al., 2005), which is a crucial ingredient for maps of large scale structure (McCleary et al., 2015; 2020; Scognamiglio, 2024).
@@ D.2 Training and Evaluating
Our models and training scripts are adapted from Pandya et al. (2025a) and https://github.com/deepskies/SIDDA. The galaxy datasets are initially sourced from https://zenodo.org/records/14583107, and there is a script to produce the datasets with PSF blurring in our artifact. We compute ECE using the same approximations as in Equations 15 - 17. We summarize the number of parameters for each model in table˜2 below:
Model CNN $C_{2}$ $C_{4}$ $C_{6}$ $C_{8}$ $C_{10}$ $C_{12}$
@@ Appendix E Chemical Property Experiment Details
We leave further compaison and model development as an opportunity for future work. Other potential baselines could include Equiformer (Liao and Smidt, 2022), EquiformerV2 (Liao et al., 2023), Graphormer (Shi et al., 2022), or Graphormer with data augmentation. The model we use in this work has $36,997,125$ parameters.